OpenStudy (anonymous):

which of the following statements are true for all values of c? I. lim f(x)=0 =>lim |f(x)|=0 x->c x->c II. lim |f(x)|=0 =>lim f(x)=0 x->c x->c are they both true? both false, or is either one true but the other false? HELP!

OpenStudy (anonymous):

got me thinking, but i would say true. not true if it wasn't zero, but since it is zero i think it is right johnny? i would love to see a counterexample

OpenStudy (anonymous):

so you would say both are true?

OpenStudy (anonymous):

oh silly me we go from the definition.

OpenStudy (anonymous):

suppose f \[\mathbb{R}-\mathbb{R}\] C and 0 are \[\epsilon \mathbb{R}\] then there exists \[\delta>0 \]

OpenStudy (anonymous):

what does it mean to say \[\lim_{x\rightarrow c}f(x)=L\]? it means given \[\epsilon > 0\] there is a \[\delta >0\] such that if \[|x-c|<\delta\] \[|f(x)-L|<\epsilon\]

OpenStudy (anonymous):

replace L by 0 and i believe you have what you need

OpenStudy (anonymous):

beat me to the punch, fast at typing haha

OpenStudy (anonymous):

ok will do.

OpenStudy (anonymous):

hmm maybe there is something to the second one though?

OpenStudy (anonymous):

no there isn't

OpenStudy (anonymous):

so 1 is true while 2 is false

OpenStudy (anonymous):

its been awhile since ive used the limit theorem, reminds me of undergrad, proving triangle inequality

OpenStudy (anonymous):

I just didn't know the rules as they applied to absolute values because we never really went over that.

OpenStudy (anonymous):

i will say they are both true. they say the same thing